A single card is drawn from a deck. Find the probability of selecting a 10 or a diamond.


Answer by HyperBrain(694)

There are 4 cards with a 10: 10 club, 10 diamond, 10 heart, 10 spade
There are 13 diamond cards:
Therefore, the number of ways to get a 10 or a diamond is (4+13)-1=16
There are 52 cards in the deck.
So, the probability is 16/52= 4/13
Power up,
HyperBrain!